OpenStudy (sloppycanada):

Write the equation of a parabola with a vertex at (-5, 2) and a directrix y = -1. y-2 = -1/12(x+5)^2

OpenStudy (sloppycanada):

\[y-2 = -\frac{ 1 }{ 12 }(x+5)^2\]

OpenStudy (amoodarya):

 \[(x-x_0)^2=4p(y-y_0)\\(x-(-5))^2=4*3(y-2)\]

OpenStudy (amoodarya):

what you write ,is